The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

Standard Hours Standard Rate Standard Cost
Motor tune-up 2.50 $31.00 $77.50

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 56 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

Labor rate variance $ 146 F
Labor spending variance $ 102 U

Required:

1. Determine the number of actual labor-hours spent on tune-ups during the week.

2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1

Labor efficiency variance = 102U - 146F = 248 U

(SH-AH) *SR = - 248

(2.50*56 - AH) *31 = - 248

(140 - AH) = - 8

AH = 148

Actual hours = 148

Labpr rate variance = (SR-AR) *AH

(31 - AR) *148 = 146

(31 - AR) = 0.99

Actual hourly rate = 30.01

Note that I have rounded some numbers. Comment if you face any issues